OpenStudy (anonymous):

(6x+3) (7x+5) How do I simplify this?

OpenStudy (anonymous):

by "simplify" you mean "multiply" yes?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

F.O.I.L. 35x^2+51x+15

OpenStudy (anonymous):

you have to do four multiplications \[(6x+3) (7x+5)=6x\times 7x+6x\times 5+3\times 7x+3\times 5\]

OpenStudy (anonymous):

then \[42x^2+30x+21x+15\] \[42x^2+51x+15\]

OpenStudy (anonymous):

yeah thats right i cant multiply ha

OpenStudy (anonymous):

Thanks the step by step was pretty useful.

OpenStudy (anonymous):

yw